Calculate the total number of ions present in 22.0 g of SrBr2.
1
Expert's answer
2018-10-04T07:20:01-0400

SrBr2 = Sr^2+ + 2Br^-1
We have 1,806 x 10^23 ions of 1 moll SrBr2
n= m/Mr. n= 22/248 n= 0.0887096774
1,806 x 10^23 x 0.0887096774= 1.60209677x 10^22